OpenStudy (anonymous):

y varies inversely with twice the value of x. When =8,x=2. Find y when =8

myininaya (myininaya):

y=k*1/(2x)

OpenStudy (anonymous):

use the same approach: x=a*y 2=a*8 a=2/8=1/4 then if y=8 x=1/4 * 8 x=2

myininaya (myininaya):

when y=8, then x=2? find y when x=8? i suppose this is what you mean to write 8=k*1/(2*2) 4*8=k 32=k so we have y=32*1/(2x) y=16/(x) y=16/8 y=2